# url]https://orcid.org/$0$-0001-7408-6925

### 26 Nov 2023 v1

@@ 3.3.4 Sentiment analysis tools
The purpose of using already trained packages and tools would be for the quick execution and implementation of sentiment calculation. Even though the performance of such techniques might not be cutting-edge, it does provide a baseline output which could lead to significant features in the dataset. Typically these consist of two types, Rule-Based and Embedding based techniques.
NLTK VADER sentiment analysis is an NLP algorithm that works with a combination of lexicon, grammatical rules, and syntactical conventions. It measures both the polarity of the text and its intensity. The lexicon dictionary used not only comprises of word sentiments but also phrases and emoticons. This would make such a technique ideal for social media text and more informal text sentiment analysis. Grammatical rules are established with heuristic rules that handle adverbs, conjunctions, punctuation capitalization, etc. An output attribute called compound presents the sum of valence scores for each word in the lexicon and shows the sentiment degree rather than value. Compound values for positive sentiment are >=0.05, neural is [-0.05, 0.05] and negative is <=0.05.
TextBlob from NLTK also utilizes a similar approach to Vader but TextBlob assesses the final sentiment of a sentence by taking a pooling operation such as the average of polarities. This, however, might have negative effect when there are an equal number of positive and negative words in the text sentence. TextBlob also provides subjectivity which quantifies the amount of factual information and subjective opinion present in the text. If the texts have greater subjectivity, personal opinion has taken the place of objective knowledge. One more key difference is that TextBlob is also suitable for non-social media text articles. Flair is another tool that uses word embedding classification for the generation of sentiment. Flair was trained on the IMBD dataset.
@@ 3.3.5 Sentiment Analysis using Sentiment Dictionaries
pos and neg is the number of positive and negative words found in the text and count is the total number of words in each sentence. However, the Loughran-McDonald dictionary is sparse, and often sentences are not allocated any polarity
Next, Harvard IV dictionary is applied in a similar approach. Harvard IV is a Dictionary with each word tagged with positive and negative attributes, in accordance to the psychological Harvard-IV dictionary as used in the General Inquirer software. Harvard IV is less sparse. These methods are extremely fast and require minimal computational power. Though there is no fine tuning they task specific for Financial news.
@@ 3.3.6 Emotion Analysis with etf-idf vectors
The tagging of emotions to each word is done using 2 different lexicons.
The first one is a conjunction of NRC Lexicon with gold standard lexicon from [45]. On thorough inspection, it was found that these two lexicons have emotion tags for the most accepted core emotions. NRC lexicon comprised of Anger, Anticipation, Disgust, Fear, Joy, Sadness, Surprise, and Trust. However, the preceding lexicon only had Joy, Anger, Sadness, Fear, and Disgust. The purpose of considering two lexicons is to maximize the number of unique words captured. NRC Lexicon comprised of 14182 and the second lexicon 13915 respectively. On analysis, it was found that 4915 words were different from the preceding lexicon which were not present in the NRC lexicon. Hence, on nulling non-existing emotion features a concatenation was performed to get a total of 19097 words and 8 different emotions in our final NRC lexicon.
@@
Preprocessing: Before the computation of etf-idf values, POS tagging is performed on each document. It is to be noted that NLTK POS tagging is context dependent hence it was executed on minimal prepossessing before stop word word removal in Preprocessing Pipeline3. After POS tagging, only adjectives, Adverbs and interjections were filtered out in each document. Nouns and verbs were first tested for emotion analysis but resulted in poor etf-idf values hence were removed. POS tag filtering helps in extracting minimal essential context from each document which is valuable for emotion analysis.
@@
Where df(emo) is the number of documents in which the respective emotion words are present. N is the number of documents. Since we calculate emotions per day this would be all the documents extracted for that specific day. One is added in the denominator to see that df values lie in the Log domain. The same formulas for etf-idf are used for NRC and senticnet with each emotion feature set. The final dataset however does not include senticnet features, only 8 emotions were included. This is because of the excessive dilutions of emotions in senticnet due to too many features.
@@ 3.4 Sentiment Analysis using Pre-Training of Language models
Because of the specialized vocabulary employed in the financial arena, general-purpose models are insufficiently effective. Added to this preserving global context, and passing context through the text article is necessary for better sentiment results.Pre-trained language models, which need fewer labeled instances and may be progressively trained on domain-specific corpora, may be able to help with this issue which we shall test in this section.
We appled transfer learning using various BERT architectures in our work. This was done by using continued pre-training over already pre-trained BERT architectures to transfer the models to perform better for our domain-specific task. Our hypothesis is that further pre-training to our extensive dataset should bring considerable performance improvement for polarity calculations. One feature of the transfer learning techniques is the capacity to continue pre-training of LM’s using domain-specific unlabeled corpora. Consequently, the model can pick up on semantic relationships. In the desired domain’s written content, which is presumably distinct from a generic corpus in terms of distribution. This strategy has particular promise for a specialized field like finance, where the terminology and language are vastly different from those in a general field.
@@ 3.4.1 Dataset
We integrate two standard financial sentiment analysis datasets for our task. Financial Phrasebank comprises of 4845 phrases in total in English that were randomly chosen from financial news articles in the LexisNexis database[50]. Financial Experts with backgrounds in business and finance annotated these sentences. The annotators were instructed to assign labels based on how they believed that the data in the phrase might impact the stock price of the listed firm. The collection also contains statistics on the degrees of agreement amongst annotators on certain sentences. We only utilize degree with 66 percent agreement. This dataset has 4217 labeled rows.
Sentiment FiQA For the WWW ’18 conference’s financial opinion mining and question-answering challenge, a dataset called FiQA [51] was produced. Only the first task that is, Aspect-based financial sentiment analysis dataset is made use of, which consists of 1,174 tweets and news stories related to finance together with the accompanying sentiment score. The labels of the dataset are regressive with continuous values between [-1,1], with 1 having the highest polarity, in contrast to Financial Phrasebank. To make both the datasets uniform for conjunction, we convert these labels into positive, negative, and neutral. On manual inspection and verification by financial experts, we find that for best results positive is [0.15,1], negative is [-0.15,-1], and neutral is the range in between. Each sample additionally specifies which financial institution is being mentioned in the statement. We don’t utilize this information.
@@
As stated in [52], to overcome Catastrophic forgetting while fine-tuning BERT-Like models; Slanted triangular learning rate, Gradual Unfreezing, and discriminative fine-tuning are employed with FinBERT, BERT, and FinancialBERT. This applies to both pre-trained and non-pre-trained models in our analysis. The rationale for gradual unfreezing and discriminative fine-tuning is that higher-level characteristics require more fine-tuning in comparison to lower-level ones since language modeling information is largely available at lower levels.
@@ 3.4.2 Evaluation Metrics
Three metrics were utilized to evaluate our sentiment classification models: macro F1 score, Area Under Curve(AUC)(One vs One) and Accuracy. It has to be noted that our respective dataset suffers from High-class imbalance problems, Positive and Neutral labels are highly dominant. One vs One AUC computes the average AUC for each pairwise combination of classes that is feasible. Disregarding the disparity in class when there is class imbalance. The area under the curve is extensively used for the performance of multi-class problems and is a recognized measure[53]. Macro F1 average computes F1 scores for every class before averaging them. This too is recommended for imbalanced datasets.
@@ 3.4.3 BERT vs Distil Architectures(RQ1)
To answer this research question we compare various set of BERT architectures. Vanilla BERT fine-tuned and FinBERT[52] (which is already pre-trained on TRC2-financial text data) fine-tuned on the labeled financial dataset respectively, DistilBERT and DistilRoBERTa fine-tuned respectively. The results obtained by finetuning the aforementioned models are shown in Table(3)
@@
DistilRoBERTa is the distilled version of RoBERTa base. The training was similar to DistilBERT. The difference is that RoBERTa sharply performs better than BERT, this was due to different hyperparameters while training. So presumably DistilRoBERTa shall perform better. The model includes 12 heads, 768 dimensions, 6 layers, and 82M parameters in total (compared to 125M parameters for RoBERTa-base). DistilRoBERTa typically performs twice as well in relation to Roberta-base asymptotically.
@@
For FinBERT in terms of variation of performance with Batch size, they perform similarly with batch size of 16 with the best result. Other hyperparameters for best fit include a learning rate of 2e-5(recommended by BERT), epochs of 4, a warm-up proportion of 0.2, and other parameters defaulting.
BERT fine-tuning underperforms FinBERT as stated in [52] even for the dataset under our consideration. However, the difference between vannila fine tuning and our work is we apply guarding techniques against catastrophic forgetting for BERT architecture too. This means that under the right conditions pre-training BERT architecture on domain-specific data does improve accuracy. BERT performs similarly with batch size variation with slightly better results for a batch size of 32. All other parameters for best fit are similar to FinBERT.
DistilRoBERTa and DistilBERT outperform FinBERT and BERT in performance. DistilBERT outperforms in terms of accuracy and F1-Macro score but lags against finBERT in Area Under Curve. The reason for Distil-based architectures to perform better to fine-tuned datasets might be because of the reduction in the number of parameters and layers resulting in better information encoding and capture. DistilRoBERTa and DistilBERT performed similarly with batch sizes of 16 and 8 respectively, we utilize batch size 16 for lesser training time and faster computation. Other hyperparameters remain similar for both models which include a learning rate of 2e-5, epochs 5, and weight decay of 0.01, and other parameters defaulting.
@@ 3.4.4 Effect of Extensive pre-training(RQ2)
FinancialBERT is a BERT model for domain-specific language representation that has been trained on huge financial corpora. It has a very similar architecture to BERT. It was trained on 1.8M TRC2 financial news articles, 400 thousand Bloomberg News articles, and a number of Corporate Reports and Earning calls. The difference between the actual FinancialBERT and our execution is we guard it against catastrophic forgetting by using Slanted triangular learning rate, Gradual Unfreezing, and Discriminative fine-tuning. FinancialBERT is available at FinancialBERT. In this section, we try to answer if extensive pre-training of BERT architecture in a financial domain positively affects sentiment classification on fine-tuning. We fine-tune the model on our labeled financial dataset. The results in comparison to other aforementioned BERT architectures are shown in Table(4). FinancialBERT performs best with a batch size of 32 with a gradual performance decline with lower batch sizes. To check for pre-training effects we use similar hyperparameters used for FinBERT fine-tuning. In other domains where NLP is used research in [55] shows that additional pretraining does improve performance.
@@ 3.4.5 Effect of pretraining on Collected Domain Data(RQ3)
Our next research question has a thorough check over performance enhancement of BERT type models by pre-training it over our collected Domain Data consisting of around the 1.5 million entries collected by us. BERT architectures come with two types of pre-training, Masked Language Modelling(MLM) pre-training and Next Sentence Prediction(NSP). We could either use only MLM pre-training or both MLM and NSP pertaining, utilizing the data at hand. This was followed by the authors of BERT. However, there would be compromises in computational usage and time complexity. It is to be noted that NSP along with MLM typically gives more enhanced results as compared to only MLM. To try all variants we train vannila BERT with MLM+NSP pretraining, FinBERT and FinancialBERT with only MLM and distilRoBERTa with its default MLM pre-training. DistilRoBERTa was orginally pre-trained using only MLM. Part of the reason for implementing only MLM over FinBERT and FinancialBERT was due to the failure of integration of these architectures with existing pre-training NSP source codes. The results and performance is shown in Table(5)
From Table(5) we note that there has been a notable improvement in the performance of FinBERT from 0.83 to 0.85. This shines a light on the efficacy of the collected data and its evaluability toward pre-training in the domain of finance. There seems to be no enhancement in performance in FiancialBERT after domain pretraining and F1 remains at 0.84, this could be because of the already extensive pre-training that was performed over financial data. It is worth to be mentioned that FinBERT outperformed FinancialBERT after pre-training over our Financial domain data when it failed to do so before(RQ2)
Further, there seems to be a notable loss in the performance of BERT pre-training using (MLM+NSP). The reason for this could be because of context irrelevance learned while Next Sentence Pre-training was performed over Domain Data. The decrement in the performance of DistilRoBERT could be because of the same reason of catastrophic forgetting while Domain MLM pre-training was performed.
@@ 3.4.6 Spearman correlation of stock returns with Sentiment(RQ4)
After collection of data which is specific to each stock and date. We created a pipeline for sentiment analysis using FinBERT pre-trained and fine-tuned on our collected corpus. The categories of data available to us are as follows; tweets, news archive snippets, Radio transcripts, and finally full news articles extracted from 4 APIs. Every data entry is annotated with the date of publishing and respective stock. Sentiment analysis is then performed for data categories separately. The reason for this is to show separate correlations and compare data quality. Further, each respective column output could be taken into a weighted average to enhance final prediction results giving Autonomy to other researchers making use of the data.
@@
Finally the pipeline results in five outputs for each date for each stock respectively. The outputs are Twitter-sentiment, Twitter-emotion, News-Archive-Sentiment, Radio-transcript-sentiment, and News-Articles(API)-sentiment. These outputs are based on different data extraction techniques that were mentioned. Twitter emotion results in 9 different emotion values for each day. Further, to show the superiority of data quality and sentiment pipeline, we calculated stock return Spearman correlation with sentiment analysis for each respective day. Stock return is the relative change in yesterday’s stock closing price w.r.t today’s closing price. The nonparametric Spearman rank-order correlation coefficient is used to evaluate the strength and direction of a monotonic connection between two data sets. It ranges from -1 to +1, where -1 and +1 indicate a precise monotonic relationship. A correlation of 0 implies no association between the datasets. Positive correlations signify that as one variable (x) increases, the other variable (y) also increases, while negative correlations indicate that as x increases, y decreases.[56]. We compute Spearman p values for each respective day with stock return.
@@
Further, we perform a similar correlation analysis with emotion analysis for each day for every stock. The correlation p-values for some emotions are given in Table(7). Emotion analysis was performed only for Twitter data. The idea is that individual investor emotions play a vital role in the macro movement of stocks. We capture 9 emotions; anger, anticipation, disgust, fear, joy, sadness, surprise, and trust. Positive emotions such as joy and trust are expected to have a positive correlation with stock return and the rest of the negative emotions are expected to have negative emotions. On correlation check, we find out that, anger, disgust, joy, and surprise have a neutral correlation with returns. From Table(7) the following observations can be made; sadness, anger, and fear have expected negative correlations. Trust, anticipation, disgust, joy, and surprise have neutral correlations. A very similar correlation trend is shown for all stocks for mentioned emotions.

### 01 Jan 2024 v3

@@ 3.3.4 Sentiment analysis tools
The purpose of using already trained packages and tools would be for the quick execution and implementation of sentiment calculation. Even though the performance of such techniques might not be cutting-edge, it does provide a baseline output, which could lead to significant features in the dataset. Typically, these consist of two types: Rule-Based and Embedding-based techniques.
NLTK VADER sentiment analysis is an NLP algorithm that works with a combination of lexicon, grammatical rules, and syntactical conventions. It measures both the polarity of the text and its intensity. The lexicon dictionary used is not only comprised of word sentiments but also phrases and emoticons. This would make such a technique ideal for social media text and more informal text sentiment analysis. Grammatical rules are established with heuristic rules that handle adverbs, conjunctions, punctuation capitalization, etc. An output attribute called compound presents the sum of valence scores for each word in the lexicon and shows the sentiment degree rather than the value. Compound values for positive sentiment are >=0.05, neural is [-0.05, 0.05] and negative is <=0.05.
TextBlob from NLTK also utilizes a similar approach to Vader but TextBlob assesses the final sentiment of a sentence by taking a pooling operation such as the average of polarities. This, however, might have a negative effect when there are an equal number of positive and negative words in the text sentence. TextBlob also provides subjectivity, which quantifies the amount of factual information and subjective opinion present in the text. If the texts have greater subjectivity, personal opinion has taken the place of objective knowledge. One more key difference is that TextBlob is also suitable for non-social media text articles. Flair is another tool that uses word embedding classification for the generation of sentiment. Flair was trained on the IMBD dataset.
@@ 3.3.5 Sentiment Analysis using Sentiment Dictionaries
pos and neg is the number of positive and negative words found in the text and count is the total number of words in each sentence. However, the Loughran-McDonald dictionary is sparse, and often sentences are not allocated any polarity.
Next, Harvard IV dictionary is applied in a similar approach. Harvard IV is a Dictionary with each word tagged with positive and negative attributes, in accordance with the psychological Harvard-IV dictionary as used in the General Inquirer software. Harvard IV is less sparse. These methods are extremely fast and require minimal computational power. Though there is no fine tuning, they are task-specific for financial news.
@@ 3.3.6 Emotion Analysis with etf-idf vectors
The tagging of emotions to each word is done using two different lexicons.
The first one is a conjunction of NRC Lexicon with gold standard lexicon from [45]. On thorough inspection, it was found that these two lexicons have emotion tags for the most accepted core emotions. NRC lexicon comprised of Anger, Anticipation, Disgust, Fear, Joy, Sadness, Surprise, and Trust. However, the preceding lexicon only had Joy, Anger, Sadness, Fear, and Disgust. The purpose of considering two lexicons is to maximize the number of unique words captured. NRC Lexicon comprised 14182 and the second lexicon 13915 respectively. Upon analysis, it was found that 4915 words were different from the preceding lexicon, which were not present in the NRC lexicon. Hence, on nulling non-existing emotion features a concatenation was performed to get a total of 19097 words and eight different emotions in our final NRC lexicon.
@@
Preprocessing: Before the computation of etf-idf values, POS tagging is performed on each document. It is to be noted that NLTK POS tagging is context dependent hence it was executed on minimal prepossessing before stop word word removal in Preprocessing Pipeline3. After POS tagging, only adjectives, Adverbs and interjections were filtered out in each document. Nouns and verbs were first tested for emotion analysis but resulted in poor etf-idf values; hence were removed. POS tag filtering helps extract minimal essential context from each document, which is valuable for emotion analysis.
@@
Where df(emo) is the number of documents in which the respective emotion words are present. N is the number of documents. Since we calculate emotions per day, this would be all the documents extracted for that specific day. One is added in the denominator to see that df values lie in the Log domain. The same formulas for etf-idf are used for NRC and senticnet with each emotion feature set. The final dataset however does not include senticnet features, only eight emotions were included. This is because of the excessive dilution of emotions in sentences due to too many features.
@@ 3.4 Sentiment Analysis using Pre-Training of Language models
Because of the specialized vocabulary employed in the financial arena, general-purpose models are insufficiently effective. Added to this preserving global context, and passing context through the text article is necessary for better sentiment results. Pre-trained language models, which need fewer labeled instances and may be progressively trained on domain-specific corpora, may be able to help with this issue, which we shall test in this section.
We applied transfer learning using various BERT architectures in our work. This was done by using continued pre-training over already pre-trained BERT architectures to transfer the models to perform better for our domain-specific task. Our hypothesis is that further pre-training to our extensive dataset should bring considerable performance improvement for polarity calculations. One feature of the transfer learning techniques is the capacity to continue pre-training of LMs using domain-specific unlabeled corpora. Consequently, the model can pick up on semantic relationships. In the desired domain’s written content, which is presumably distinct from a generic corpus in terms of distribution. This strategy has particular promise for a specialized field like finance, where the terminology and language are vastly different from those in a general field.
@@ 3.4.1 Dataset
We integrate two standard financial sentiment analysis datasets for our task. Financial Phrasebank comprises 4845 phrases in total in English that were randomly chosen from financial news articles in the LexisNexis database[50]. Financial Experts with backgrounds in business and finance annotated these sentences. The annotators were instructed to assign labels based on how they believed that the data in the phrase might impact the stock price of the listed firm. The collection also contains statistics on the degrees of agreement amongst annotators on certain sentences. We only utilize degrees with 66 percent agreement. This dataset has 4217 labeled rows.
Sentiment FiQA For the WWW ’18 conference’s financial opinion mining and question-answering challenge, a dataset called FiQA [51] was produced. Only the first task that is, the Aspect-based financial sentiment analysis dataset is made use of, which consists of 1,174 tweets and news stories related to finance together with the accompanying sentiment score. The labels of the dataset are regressive with continuous values between [-1,1], with 1 having the highest polarity, in contrast to Financial Phrasebank. To make both the datasets uniform for conjunction, we convert these labels into positive, negative, and neutral. On manual inspection and verification by financial experts, we find that for best results positive is [0.15,1], negative is [-0.15,-1], and neutral is the range in between. Each sample additionally specifies which financial institution is being mentioned in the statement. We don’t utilize this information.
@@
As stated in [52], to overcome Catastrophic forgetting while fine-tuning BERT-Like models, Slanted triangular learning rate, Gradual Unfreezing, and discriminative fine-tuning are employed with FinBERT, BERT, and FinancialBERT. This applies to both pre-trained and non-pre-trained models in our analysis. The rationale for gradual unfreezing and discriminative fine-tuning is that higher-level characteristics require more fine-tuning in comparison to lower-level ones since language modeling information is largely available at lower levels.
@@ 3.4.2 Evaluation Metrics
Three metrics were utilized to evaluate our sentiment classification models: macro F1 score, Area Under Curve(AUC)(One vs One), and Accuracy. It has to be noted that our respective dataset suffers from high-class imbalance problems, and positive and neutral labels are highly dominant. One vs. One AUC computes the average AUC for each pairwise combination of classes that is feasible. Disregarding the disparity in class when there is class imbalance. The area under the curve is extensively used for the performance of multi-class problems and is a recognized measure[53]. Macro F1 average computes F1 scores for every class before averaging them. This too is recommended for imbalanced datasets.
@@ 3.4.3 BERT vs Distil Architectures(RQ1)
To answer this research question, we compare various set of BERT architectures. Vanilla BERT fine-tuned and FinBERT[52] (which is already pre-trained on TRC2-financial text data) fine-tuned on the labeled financial dataset respectively, DistilBERT and DistilRoBERTa fine-tuned respectively. The results obtained by finetuning the aforementioned models are shown in Table(3)
@@
DistilRoBERTa is the distilled version of the RoBERTa base. The training was similar to DistilBERT. The difference is that RoBERTa sharply performs better than BERT, this was due to different hyperparameters while training. So, presumably, DistilRoBERTa shall perform better. The model includes 12 heads, 768 dimensions, six layers, and 82M parameters in total (compared to 125M parameters for RoBERTa-base). DistilRoBERTa typically performs twice as well in relation to Roberta-base asymptotically.
@@
For FinBERT in terms of variation of performance with Batch size, they perform similarly with a batch size of 16 with the best result. Other hyperparameters for best fit include a learning rate of 2e-5(recommended by BERT), epochs of 4, a warm-up proportion of 0.2, and other parameters defaulting.
BERT fine-tuning underperforms FinBERT as stated in [52] even for the dataset under our consideration. However, the difference between Vanilla fine-tuning and our work is we apply guarding techniques against catastrophic forgetting for BERT architecture, too. This means that under the right conditions, pre-training BERT architecture on domain-specific data does improve accuracy. BERT performs similarly with batch size variation with slightly better results for a batch size of 32. All other parameters for best fit are similar to FinBERT.
DistilRoBERTa and DistilBERT outperform FinBERT and BERT in performance. DistilBERT outperforms in terms of accuracy and F1-Macro score but lags against finBERT in Area Under Curve. The reason why Distil-based architectures tend to perform better in fine-tuning datasets might be because of the reduction in the number of parameters and layers, resulting in better information encoding and capture. DistilRoBERTa and DistilBERT performed similarly with batch sizes of 16 and 8 respectively, we utilize batch size 16 for lesser training time and faster computation. Other hyperparameters remain similar for both models which include a learning rate of 2e-5, epochs 5, a weight decay of 0.01, and other parameters defaulting.
@@ 3.4.4 Effect of Extensive pre-training(RQ2)
FinancialBERT is a BERT model for domain-specific language representation that has been trained on huge financial corporations. It has a very similar architecture to BERT. It was trained on 1.8M TRC2 financial news articles, 400 thousand Bloomberg News articles, and a number of corporate reports and earnings calls. The difference between the actual FinancialBERT and our execution is we guard it against catastrophic forgetting by using a Slanted triangular learning rate, Gradual Unfreezing, and Discriminative fine-tuning. FinancialBERT is available at FinancialBERT. In this section, we try to answer if extensive pre-training of BERT architecture in a financial domain positively affects sentiment classification on fine-tuning. We fine-tune the model on our labeled financial dataset. The results in comparison to other aforementioned BERT architectures are shown in Table(4). FinancialBERT performs best with a batch size of 32 with a gradual performance decline with lower batch sizes. To check for pre-training effects, we use similar hyperparameters to those used for FinBERT fine-tuning. In other domains where NLP is used research in [55] shows that additional pretraining does improve performance.
@@ 3.4.5 Effect of pretraining on Collected Domain Data(RQ3)
Our next research question thoroughly checks the performance enhancement of BERT type models by pre-training it over our collected Domain Data, consisting of around the 1.5 million entries collected by us. BERT architectures come with two types of pre-training: Masked Language Modelling(MLM) pre-training and Next Sentence Prediction(NSP). We could either use only MLM pre-training or both MLM and NSP pertaining to utilizing the data at hand. This was followed by the authors of BERT. However, there would be compromises in computational usage and time complexity. It is to be noted that NSP along with MLM typically gives more enhanced results as compared to only MLM. To try all variants, we train vanilla BERT with MLM+NSP pretraining, FinBERT, and FinancialBERT with only MLM and distilRoBERTa with its default MLM pre-training. DistilRoBERTa was originally pre-trained using only MLM. Part of the reason for implementing only MLM over FinBERT and FinancialBERT was due to the failure of integration of these architectures with existing pre-training NSP source codes. The results and performance is shown in Table(5)
From Table(5) we note that there has been a notable improvement in the performance of FinBERT from 0.83 to 0.85. This shines a light on the efficacy of the collected data and its evaluability toward pre-training in the domain of finance. There seems to be no enhancement in performance in FiancialBERT after domain pretraining and F1 remains at 0.84, this could be because of the already extensive pre-training that was performed over financial data. It is worth mentioning that FinBERT outperformed FinancialBERT after pre-training over our Financial domain data when it failed to do so before(RQ2)
Further, there seems to be a notable loss in the performance of BERT pre-training using (MLM+NSP). The reason for this could be because of context irrelevance learned while Next Sentence Pre-training was performed over Domain Data. The decrement in the performance of DistilRoBERT could be because of the same reason for catastrophic forgetting while Domain MLM pre-training was performed.
@@ 3.4.6 Spearman correlation of stock returns with Sentiment(RQ4)
After collection of data which is specific to each stock and date. We created a pipeline for sentiment analysis using FinBERT pre-trained and fine-tuned on our collected corpus. The categories of data available to us are as follows: tweets, news archive snippets, radio transcripts, and full news articles extracted from 4 APIs. Every data entry is annotated with the date of publishing and respective stock. Sentiment analysis is then performed for data categories separately. The reason for this is to show separate correlations and compare data quality. Further, each respective column output could be taken into a weighted average to enhance final prediction results, giving Autonomy to other researchers making use of the data.
@@
Finally the pipeline results in five outputs for each date for each stock respectively. The outputs are Twitter-sentiment, Twitter-emotion, News-Archive-Sentiment, Radio-transcript-sentiment, and News-Articles(API)-sentiment. These outputs are based on different data extraction techniques that were mentioned. Twitter emotion results in 9 different emotion values for each day. Further, to show the superiority of data quality and sentiment pipeline, we calculated stock return Spearman correlation with sentiment analysis for each respective day. Stock return is the relative change in yesterday’s stock closing price w.r.t today’s closing price. The non-parametric Spearman rank-order correlation coefficient is used to evaluate the strength and direction of a monotonic connection between two data sets. It ranges from -1 to +1, where -1 and +1 indicate a precise monotonic relationship. A correlation of 0 implies no association between the datasets. Positive correlations signify that as one variable (x) increases, the other variable (y) also increases, while negative correlations indicate that as x increases, y decreases.[56]. We compute Spearman p values for each respective day with stock return.
@@
Further, we perform a similar correlation analysis with emotion analysis for each day for every stock. The correlation p-values for some emotions are given in Table(7). Emotion analysis was performed only for Twitter data. The idea is that individual investor emotions play a vital role in the macro movement of stocks. We capture nine emotions: anger, anticipation, disgust, fear, joy, sadness, surprise, and trust. Positive emotions such as joy and trust are expected to have a positive correlation with stock return and the rest of the negative emotions are expected to have negative emotions. On correlation check, we find out that, anger, disgust, joy, and surprise have a neutral correlation with returns. From Table(7) the following observations can be made: sadness, anger, and fear have expected negative correlations. Trust, anticipation, disgust, joy, and surprise have neutral correlations. A very similar correlation trend is shown for all stocks for mentioned emotions.